OpenStudy (anonymous):

Cobalt-56 has a decay constant of 8.77 × 10^-3 (which is equivalent to a half life of 79 days). How many days will it take for a sample of cobalt-56 to decay to 62% of its original value? A) 1.9 days B) 23.7 days C)54.5 days D)79 days E) 141 days

OpenStudy (aaronq):

Use first-order integrated rate law: \(\sf \large ln[A]_t=-kt+ln[A]_o\) where \(\sf ln[A]_o\) is the initial concentration of the substance \(\sf k\) is the decay constant \(\sf t\) is time \(\sf ln[A]_t\) is the concentration of substance at time \(\sf t\)

OpenStudy (anonymous):

The function of the half life is:\[A=A_o\cdot e^{-\lambda\cdot t}\] \[A=100\%\cdot e^{-8.77\times10^{-3}\cdot t}\]Now we have to replace A by 62%\[62\%=100\%\cdot e^{-8.77\times10^{-3}\cdot t}\]\[0.62=e^{-8.77\times10^{-3}\cdot t}\]\[ln(0.62)=(-8.77\times10^{-3}\cdot t)\cdot ln(e)\]\[\boxed{t\approx54.5}\]

OpenStudy (anonymous):

The answer is C

OpenStudy (anonymous):

give me a medal for liking cage free chicken! <3 plus, i don't have any yet.